OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 25?

OpenStudy (anonymous):

which function is below?

OpenStudy (ranga):

Steps: Write the function as y = x^2 - 25 Solve for x in terms of y. Interchange x and y.

OpenStudy (anonymous):

\[y=x^2 -25\] \[x=y^2 -25\] \[\sqrt{x+25}=\sqrt{y^2}\] \[y=\pm \sqrt{x+25}\]